OpenStudy (anonymous):

Indicate which property is illustrated in Step 5. Step 1 7x + 3 + 11x + 8 = (7x + 3) + (11x + 8) Step 2 = 7x + (3 + 11x) + 8 Step 3 = 7x + (11x + 3) + 8 Step 4 = (7x + 11x) + (3 + 8) Step 5 = (7 + 11)x + (3 + 8) Step 6 = 18x + 11

OpenStudy (anonymous):

@MAli13chineta @kl0723 @KatClaire

OpenStudy (kl0723):

it factored out the x

OpenStudy (anonymous):

that be the distributive law

OpenStudy (kl0723):

but then it goes back to the same things :P

OpenStudy (anonymous):

\[\huge 7x+11x=(7+11)x\] \[\huge ba+ca=(b+c)a\]